OpenStudy (anonymous):

how do yo find the substitution method for 5x-3y=5 and 10x+6y=20? 2x-3y=2 and 4x-6=9? 2x+4y=12 and 6x+12y=36?

OpenStudy (lilg132):

you dont click best response when i answer so i refuse to answer =)

OpenStudy (anonymous):

i just did, so answer the question, please!

OpenStudy (anonymous):

Substitution means solving one equation for x or y then plugging it into another equation. Though I would use the elimination method for both of these systems. For the first system, multiply (5x+3y=25) by -2 and you will get -10x+-6y=-50. Then u can cancel out the 10x and -10x. In the second system, multiply -3

OpenStudy (lilg132):

what about the previous question i answered !

OpenStudy (anonymous):

hey im gonna delete my answer now since u gave the best response to someone else.... jk :)

OpenStudy (anonymous):

i dont hab alot time so can you guys just help me........

OpenStudy (anonymous):

i did. look at my answer, does it help or do u need more clarification?

OpenStudy (anonymous):

uhm i need more clarification, like step by step....

OpenStudy (anonymous):

ok eq 1) (5x+3y=25) * -2= -10x + -6y =-50 eq 2) 10x + 6y = 20 the 10x and the -10x cancel out to zero. as well as the -6y and the 6y.

OpenStudy (lilg132):

baisically you want the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for the first system the answer is (-1, 3/10)